OpenStudy (devincoolage1001):

Solve for x: (3/4) x + 5/8 = 4x Plz help

OpenStudy (freckles):

you could start by getting the x terms together by subtracting (3/4)x on both sides or you could start by getting rid of the fractions by multiplying both sides by 8

OpenStudy (guoalice8):

5/8=3.75x

OpenStudy (robtobey2):

\[\text{Expand}\left[8 \left(\frac{3 x}{4}+\frac{5}{8}\right)=8 (4 x)\right] \]5 + 6 x = 32 x 26x = 5 x = 5/26

OpenStudy (devincoolage1001):

thankyou